# A Posteriori Single- and Multi-Goal Error Control and Adaptivity for Partial Differential Equations

### v1

@@ Abstract
This work reviews goal-oriented a posteriori error control, adaptivity and solver control for finite element approximations to boundary and initial-boundary value problems for stationary and non-stationary partial differential equations, respectively. In particular, coupled field problems with different physics may require simultaneously the accurate evaluation of several quantities of interest, which is achieved with multi-goal oriented error control. Sensitivity measures are obtained by solving an adjoint problem. Error localization is achieved with the help of a partition-of-unity. We also review and extend theoretical results for efficiency and reliability by employing a saturation assumption. The resulting adaptive algorithms allow to balance discretization and non-linear iteration errors, and are demonstrated for four applications: Poisson’s problem, non-linear elliptic boundary value problems, stationary incompressible Navier-Stokes equations, and regularized parabolic $p$-Laplace initial-boundary value problems. Therein, different finite element discretizations in two different software libraries are utilized, which are partially accompanied with open-source implementations on GitHub.
@@ 1 Introduction
This work is devoted to a goal-oriented a posteriori error control of single and multiple quantities of interest (QoI) and Adaptive Finite Element Methods (AFEM) for solving stationary and non-stationary, linear and non-linear partial differential equations (PDEs) and systems of PDEs. AFEM were certainly inspired by the pioneer work [12] of Babuška and Rheinboldt. Further important studies on a posteriori error controlled adaptive finite element methods are [80, 60, 158, 42, 202] to name a few. The first comprehensive monograph [213] by Verfürth was another AFEM milestone. We refer to the survey papers [98], [34], [62], and to the monographs [4], [14], [17], [116], [161], [183], [99], [184] for an overview of a posteriori error estimates and adaptive finite element techniques.
The governing discretization in this work is the finite element method [52, 50, 66, 127, 57, 100], but isogeometric analysis could be used in a similar fashion [129, 68, 35]. We consider stationary settings as well as time-dependent cases, modeled as space-time problems and solved by space-time discretization methods; see, e.g., [131, 128, 210, 200, 163, 201, 139, 205, 140] and the references therein.
@@
Conceptional developments on goal-oriented error estimation include a safe-guarded method established in [165] and guaranteed bounds derived in [5]. Abstract analyses and reformulations were presented in [103, 133]. First convergence rates of goal-oriented error estimation were obtained in [157], weighted marking [27], bounding techniques for goal-oriented error estimates for elliptic problems [138], goal oriented flux reconstruction [159], goal-oriented error estimation for the fractional-step-theta scheme [149], a partition-of-unity localization including effectivity estimates for single goal functionals [189], partition-of-unity localization for multiple goal functionals [96], space-time partition-of-unity localization [90, 209], equilibrated flux [51], linearization errors [114]. Theoretical work showing optimality of adaptive algorithms was summarized in four axioms of adaptivity in [58] and dates back to residual-based a posteriori error estimates in [80]. First rigorous convergence results for goal-oriented estimators go back to [157, 27], then [103, 126, 125], and recent findings are [30, 26, 28]. We also mention our own prior work on such theoretical advancements that include effective estimates with common upper bounds of the dual-weighted residual estimator and the error indicators [189]. Efficiency and reliability estimates for the dual-weighted residual method using a saturation assumption were shown in [94], and smart algorithms switching between solving high-order adjoint problems or using interpolations were derived in [95].
@@
Single- and multigoal-oriented error control and adaptivity has found numerous applications. These include Poisson’s problem [33, 32, 171, 96], stationary linear elasticity [176], space-time adaptivity for parabolic equations [195], space-time elasticity and the wave equation [173, 15], multi-rate discretizations of coupled parabolic/hyperbolic problems [199], $p$-Laplace problems [93, 95, 90, 181], elasto-plasticity [177, 178, 179] and visco-plasticity [146], incompressible flow [29, 49, 36, 37, 91, 94, 95, 192], transport proplems [137], transport problems with coupled flow [21, 54, 55], Boussinesq equations and coupled flow with temperature [40, 39, 87], reactive flows [25, 185, 48], uncertain inputs [38], Maxwell’s equations [130], elliptic eigenvalue problems [120, 67], modeling errors [166, 47, 174], applications to polygonal meshes [216], conforming and nonconforming approximations with inexact solvers [145], anisotropic mesh refinement [186], heterogeneous multiscale problems [65, 143, 144, 141, 79], fluid-structure interaction [115, 187, 82, 83, 84, 217, 218, 105, 212, 188, 102, 3], phase-field slits and fracture [219, 220, 221], sea ice simulations [146], optimal control [31, 24, 151, 147, 123, 214, 124, 180, 122, 89], obstacle/contact problems and variational inequalities [44, 45, 43, 196, 175, 204, 220], finite cell methods [203, 76], balancing discretization and iteration errors [148, 182, 181, 93, 77, 78], financial mathematics [111], goal-oriented model order reduction [153, 106, 107, 132, 61], neural network enhanced dual-weighted residual technologies [53, 156, 191], adaptive multiscale predictive modeling [167], and open-source software developments for goal-oriented error estimation [112] [^5] https://github.com/winnifried/dopelib, [136] [^6] https://github.com/dtm-project/dwr-diffusion, [209, 206] [^7] https://github.com/jpthiele/pu-dwr-diffusion, [^8] https://github.com/jpthiele/pu-dwr-combustion.
@@
The outline of this work is as follows. In Section 2, the notation and abstract setting are introduced. Then, in Section 3, single goal-oriented error control is discussed. Section 4 focuses on error estimation for non-standard discretizations. Next, in Section 5, multigoal-oriented error estimators are explained. In Section 6, three applications are considered, and computationally analyzed. The work is concluded in Section 7, and some current research directions are outlined.
@@ 2 Notation, Abstract Setting, Finite Element Discretization
Throughout this work, let $d$ be the space dimension. We denote by $Q:=\Omega\times I\subset\mathbb{R}^{d+1}$ the space-time domain (cylinder), where $\Omega\subset\mathbb{R}^{d}$ is the bounded and Lipschitz spatial domain with the boundary $\Gamma=\partial\Omega$, and $I:=(0,T)$ denotes the temporal domain (time interval). Here, $T>0$ is the end time value. Furthermore, we use the standard notations for Lebesgue, Sobolev, and Bochner spaces like $L_{p}(\Omega)$, $W_{p}^{k}(\Omega),\mathring{W}_{p}^{k}(\Omega),H^{k}(\Omega)=W_{2}^{k}(\Omega),H^{k}_{0}(\Omega)=\mathring{W}_{2}^{k}(\Omega)$, $L_{p}(0,T;\mathring{W}_{p}^{1}(\Omega))$ etc.; see, e.g., [50, 52, 222].
@@ 2.2.2 Finite Elements on Hypercubes $Q_{k}$
To facilitate adaptive mesh refinement and to avoid connecting elements, we use the concept of hanging nodes. Elements are allowed to have nodes that lie on the midpoints of the faces or edges of neighboring cells. In our implementations, at most, one hanging node is allowed on each face or edge. In three dimensions, this concept is generalized to subplanes and faces because we must deal with two types of lower manifolds. To enforce global continuity (i.e., global conformity), the degrees of freedom located on the interface between different refinement levels have to satisfy additional constraints. They are determined by interpolation of neighboring degrees of freedom. Therefore, hanging nodes do not carry any degrees of freedom. For more details on this, we refer to [57].
@@ 3.2 Adjoint Problem
In order to realize goal-oriented error control in this work, we use the DWR method, which requires solving an adjoint problem. This adjoint formulation follows from the Lagrange formalism and is given by: Find $z\in V$ such that
@@ 3.3 An Error Identity
$\mathcal{L}^{\prime}(\tilde{x})(x-\tilde{x})=$ $\underbrace{J^{\prime}(\tilde{u})(e)-\mathcal{A}^{\prime}(\tilde{u})(e,\tilde{z})}_{=\rho^{*}(\tilde{u},\tilde{z})(u-\tilde{u})}\underbrace{-A(\tilde{u})(e^{*})}_{=\rho(\tilde{u})(z-\tilde{z})},$
@@ 4.2 Estimating the Consistency Error
$\rho_{h}(u)(v)=-A_{h}(u)(v),\quad\rho_{h}^{*}(u,z)(v)=J^{\prime}(u)(v)-A^{\prime}(u)(v,z).$
@@ 4.2.1 Consistency Error in Galerkin Time-Stepping Methods
$A_{h}(u_{h})(v_{h})=\left(u_{h}(t_{0})-u_{0}\right)\cdot v_{h}(t_{0})+\sum_{n=1}^{N}\left(u_{h}(t_{n})-u_{h}(t_{n-1})+\Delta tf\left(t_{n},u(t_{n})\right)\right)\cdot v_{h}(t_{n}).$ (74)
@@
$A(u_{h})(v_{h})=\left(u_{h}(t_{0})-u_{0}\right)\cdot v_{h}(t_{0})\\ +\sum_{n=1}^{N}\left\{\left(u_{h}(t_{n})-u_{h}(t_{n-1}\right)\cdot v_{h}(t_{n})+\int_{t_{n-1}}^{t_{n}}f(t,u_{h}(t))\cdot v_{h}(t)\right\}.$ (76)
@@
The consistency error can be introduced as $\mathcal{S}_{h}(u)(z)=A_{h}(u)(z)-A(u)(z)$ and for the Euler method we obtain

### v2

@@ Abstract
This chapter reviews goal-oriented a posteriori error control, adaptivity and solver control for finite element approximations to boundary and initial-boundary value problems for stationary and non-stationary partial differential equations, respectively. In particular, coupled field problems with different physics may require simultaneously the accurate evaluation of several quantities of interest, which is achieved with multi-goal oriented error control. Sensitivity measures are obtained by solving an adjoint problem. Error localization is achieved with the help of a partition-of-unity. We also review and extend theoretical results for efficiency and reliability by employing a saturation assumption. The resulting adaptive algorithms allow to balance discretization and non-linear iteration errors, and are demonstrated for four applications: Poisson’s problem, non-linear elliptic boundary value problems, stationary incompressible Navier-Stokes equations, and regularized parabolic $p$-Laplace initial-boundary value problems. Therein, different finite element discretizations in two different software libraries are utilized, which are partially accompanied with open-source implementations on GitHub.
@@ 1 Introduction
This chapter is devoted to a goal-oriented a posteriori error control of single and multiple quantities of interest (QoI) and Adaptive Finite Element Methods (AFEM) for solving stationary and non-stationary, linear and non-linear partial differential equations (PDEs) and systems of PDEs. AFEM were certainly inspired by the pioneer work [13] of Babuška and Rheinboldt. Further important studies on a posteriori error controlled adaptive finite element methods are [85, 65, 167, 43, 211] to name a few. The first comprehensive monograph [223] by Verfürth was another AFEM milestone. We refer to the survey papers [105], [35], [67], and to the monographs [4], [15], [18], [125], [170], [192], [106], [193] for an overview of a posteriori error estimates and adaptive finite element techniques.
The governing discretization in this chapter is the finite element method [54, 51, 71, 136, 62, 107], but isogeometric analysis could be used in a similar fashion [138, 73, 36]. We consider stationary settings as well as time-dependent cases, modeled as space-time problems and solved by space-time discretization methods; see, e.g., [140, 137, 220, 209, 172, 210, 148, 215, 149] and the references therein.
@@
Conceptional developments on goal-oriented error estimation include a safe-guarded method established in [174] and guaranteed bounds derived in [5]. Abstract analyses and reformulations were presented in [110, 142]. First convergence rates of goal-oriented error estimation were obtained in [166], weighted marking [28], bounding techniques for goal-oriented error estimates for elliptic problems [147], goal oriented flux reconstruction [168], goal-oriented error estimation for the fractional-step-theta scheme [158], a partition-of-unity localization including effectivity estimates for single goal functionals [198], partition-of-unity localization for multiple goal functionals [103], space-time partition-of-unity localization [96, 219, 102], equilibrated flux [52], linearization errors [122]. Theoretical work showing optimality of adaptive algorithms was summarized in four axioms of adaptivity in [63] and dates back to residual-based a posteriori error estimates in [85]. First rigorous convergence results for goal-oriented estimators go back to [166, 28], then [110, 135, 134], and recent findings are [31, 27, 29]. We also mention our own prior work on such theoretical advancements that include effective estimates with common upper bounds of the dual-weighted residual estimator and the error indicators [198]. Efficiency and reliability estimates for the dual-weighted residual method using a saturation assumption were shown in [100], and smart algorithms switching between solving high-order adjoint problems or using interpolations were derived in [101].
@@
Single- and multigoal-oriented error control and adaptivity has found numerous applications. These include Poisson’s problem [34, 33, 180, 103, 61], stationary linear elasticity [185], nonlinear finite deformation elasticity [121], space-time adaptivity for parabolic equations [204], space-time elasticity and the wave equation [182, 16], multi-rate discretizations of coupled parabolic/hyperbolic problems [208], $p$-Laplace problems [99, 101, 96, 190], elasto-plasticity [186, 187, 188] and visco-plasticity [155], incompressible flow [30, 50, 37, 38, 97, 100, 101, 201], transport problems [146], transport problems with coupled flow [22, 57, 58], Boussinesq equations and coupled flow with temperature [41, 40, 93], reactive flows [26, 194, 49], biomechanics [90, 60], uncertain inputs [39], Maxwell’s equations [139], elliptic eigenvalue problems [129, 72], modeling errors [175, 48, 183], applications to polygonal meshes [226], conforming and nonconforming approximations with inexact solvers [154], anisotropic mesh refinement [195], heterogeneous multiscale problems [70, 152, 153, 150, 84], fluid-structure interaction [123, 196, 87, 88, 89, 227, 228, 112, 222, 197, 109, 3], phase-field slits and fracture [229, 230, 231], sea ice simulations [155], optimal control [32, 25, 160, 156, 132, 224, 133, 189, 131, 95], obstacle/contact problems and variational inequalities [45, 46, 44, 205, 184, 214, 230], finite cell methods [213, 81], balancing discretization and iteration errors [157, 191, 190, 99, 82, 83], financial mathematics [118], goal-oriented model order reduction [162, 113, 114, 141, 66], neural network enhanced dual-weighted residual technologies [56, 165, 200], adaptive multiscale predictive modeling [176], and open-source software developments for goal-oriented error estimation [119] [^5] https://github.com/winnifried/dopelib, [61] [^6] https://github.com/jhale/fenicsx-error-estimation, [90] [^7] https://doi.org/10.6084/m9.figshare.8128178.v1, [145] [^8] https://github.com/dtm-project/dwr-diffusion, [219, 216] [^9] https://github.com/jpthiele/pu-dwr-diffusion, [^10] https://github.com/jpthiele/pu-dwr-combustion.
@@
The outline of this book chapter is as follows. In Section 2, the notation and abstract setting are introduced. Then, in Section 3, single goal-oriented error control is discussed. Section 4 focuses on error estimation for non-standard discretizations. Next, in Section 5, multigoal-oriented error estimators are explained. In Section 6, three applications are considered, and computationally analyzed. The book chapter is concluded in Section 7, and some current research directions are outlined.
@@ 2 Notation, Abstract Setting, Finite Element Discretization
Throughout this book chapter, let $d$ be the space dimension. We denote by $Q:=\Omega\times I\subset\mathbb{R}^{d+1}$ the space-time domain (cylinder), where $\Omega\subset\mathbb{R}^{d}$ is the bounded and Lipschitz spatial domain with the boundary $\Gamma=\partial\Omega$, and $I:=(0,T)$ denotes the temporal domain (time interval). Here, $T>0$ is the end time value. Furthermore, we use the standard notations for Lebesgue, Sobolev, and Bochner spaces like $L_{p}(\Omega)$, $W_{p}^{k}(\Omega),\mathring{W}_{p}^{k}(\Omega),H^{k}(\Omega)=W_{2}^{k}(\Omega),H^{k}_{0}(\Omega)=\mathring{W}_{2}^{k}(\Omega)$, $L_{p}(0,T;\mathring{W}_{p}^{1}(\Omega))$ etc.; see, e.g., [51, 54, 232].
@@ 2.2.2 Finite Elements on Hypercubes $Q_{k}$
To facilitate adaptive mesh refinement and to avoid connecting elements, we use the concept of hanging nodes. Elements are allowed to have nodes that lie on the midpoints of the faces or edges of neighboring cells. In our implementations, at most, one hanging node is allowed on each face or edge. In three dimensions, this concept is generalized to subplanes and faces because we must deal with two types of lower manifolds. To enforce global continuity (i.e., global conformity), the degrees of freedom located on the interface between different refinement levels have to satisfy additional constraints. They are determined by interpolation of neighboring degrees of freedom. Therefore, hanging nodes do not carry any degrees of freedom. For more details on this, we refer to [62]. To ensure the approximation properties of the finite element spaces, additional conditions on the geometry of the cells are required. The two classical assumptions from the literature, e.g., [54, 51] are the so-called uniformity and the weaker concept of quasi-uniformity.
@@ 3.2 Adjoint Problem
In order to realize goal-oriented error control in this book chapter, we use the DWR method, which requires solving an adjoint problem. This adjoint formulation follows from the Lagrange formalism and is given by: Find $z\in V$ such that
@@ 3.3 An Error Identity
$\mathcal{L}^{\prime}(\tilde{x})(x-\tilde{x})=$ $\underbrace{J^{\prime}(\tilde{u})(e)-\mathcal{A}^{\prime}(\tilde{u})(e,\tilde{z})}_{=\rho^{*}(\tilde{u},\tilde{z})(u-\tilde{u})}\underbrace{-\mathcal{A}(\tilde{u})(e^{*})}_{=\rho(\tilde{u})(z-\tilde{z})},$
@@ 3.10.2 Details on Using $\tilde{u}_{h}^{(2)}$ for the Localization
In theory, there are no specific assumptions on the mesh except the usual requirements in terms of the (adaptive) finite element method like shape-regularity of the mesh; see Section 2.2. However, to make the discretization fully computable, so-called variational crimes are permitted; see, e.g., Chapter 8 in [55], For instance, one approximates the boundary $\partial\Omega$ of the spatial domain $\Omega$ or/and interfaces appropriately, and one uses numerical integration for computing the stiffness matrices, the right-hand sides, and the finite element approximation to the non-linear operator $\mathcal{A}(\cdot)$ and its computation at a finite element function / vector. This in turn depends on the mesh, and leads to additional contributions to the oscillation error. Furthermore, we have to assume that this error is sufficiently small such that the higher-order parts of the error decay fast. However, on very coarse meshes, this part can influence the discretization error estimator. We also refer to Section 4 for the analysis of non-standard discretizations.
Remark 4.
@@ 4.2 Estimating the Consistency Error
$\rho_{h}(u)(v)=-\mathcal{A}_{h}(u)(v),\quad\rho_{h}^{*}(u,z)(v)=J^{\prime}(u)(v)-\mathcal{A}^{\prime}(u)(v,z).$
@@ 4.2.1 Consistency Error in Galerkin Time-Stepping Methods
$\mathcal{A}_{h}(u_{h})(v_{h})=\left(u_{h}(t_{0})-u_{0}\right)\cdot v_{h}(t_{0})+\sum_{n=1}^{N}\left(u_{h}(t_{n})-u_{h}(t_{n-1})+\Delta tf\left(t_{n},u(t_{n})\right)\right)\cdot v_{h}(t_{n}).$ (74)
@@
$\mathcal{A}(u_{h})(v_{h})=\left(u_{h}(t_{0})-u_{0}\right)\cdot v_{h}(t_{0})+\sum_{n=1}^{N}\left\{\left(u_{h}(t_{n})-u_{h}(t_{n-1}\right)\cdot v_{h}(t_{n})+\int_{t_{n-1}}^{t_{n}}f(t,u_{h}(t))\cdot v_{h}(t)\right\}.$ (76)
@@
The consistency error can be introduced as $\mathcal{S}_{h}(u)(z)=\mathcal{A}_{h}(u)(z)-\mathcal{A}(u)(z)$ and for the Euler method we obtain
@@ 4.3.1 Non-Conforming Finite Elements
4.3.2 Non-Conforming Boundary Approximation with Nitsche’s Method
As a typical example of a non-conforming discretization, we consider the problem: Find $u\in U:=H^{1}_{0}(\Omega)$ such that
$-\Delta u+u^{2}=f\text{ in }\Omega,$ (82)
which we approximate using Nitsche’s method [171] to weakly enforce the boundary conditions, i.e. finding $u_{h}\in U_{h}$ where $U_{h}\subset H^{1}(\Omega)$ but not necessarily $u_{h}=0$ on $\partial\Omega$ such that in the weak form, we have
$A_{h}(u_{h})(v_{h}):=(\nabla u_{h},\nabla v_{h})+(u_{h}^{2},v_{h})-\langle\partial_{n}u_{h},v_{h}\rangle_{\partial\Omega}+\frac{\gamma}{h}\langle u_{h},v_{h}\rangle_{\partial\Omega}-(f,v_{h})\stackrel{{\scriptstyle!}}{{=}}0\quad\forall v_{h}\in U_{h}(\Omega),$ (83)
where $\gamma>0$ is a parameter controlling the penalization of $u_{h}=0$. The linearized adjoint to (82) is: Find $z\in U:=H^{1}_{0}(\Omega)$ such that
$-\Delta z+2uz=j^{\prime}(u)\text{ in }\Omega,$ (84)
where $j^{\prime}(u)$ is the strong form of the goal functional, i.e. $J^{\prime}(u)(v)=\langle j^{\prime}(u),v\rangle$, and which is itself discretized using Nitsche’s method in the discrete variational form
$A_{h}^{*}(u_{h})(v_{h},z_{h}):=(\nabla v_{h},\nabla z_{h})+(2u_{h}v_{h},z_{h})-\langle\partial_{n}z_{h},v_{h}\rangle_{\partial\Omega}+\frac{\gamma}{h}\langle z_{h},v_{h}\rangle_{\partial\Omega}\stackrel{{\scriptstyle!}}{{=}}J^{\prime}(u_{h})(v_{h}).$ (85)
We note that it does not hold $A^{\prime}_{h}(u)(v,z)=A^{*}_{h}(u)(v,z)$ as the primal Nitsche form (83) does not contain a symmetry term on the boundary. Here, discretizing with Nitsche and taking the adjoint does not commute.
Since $u=v=z=0$ on $\partial\Omega$ for $u,v,z\in U:=H^{1}_{0}(\Omega)$, the approach is a consistent method and the solutions $u,z\in U$ satisfy $A_{h}(u,v)=0$ and $A^{*}_{h}(u)(v,z)=J^{\prime}(u)(v)$ for $v\in U$. In the following, we will mostly work with the abstract formulations $A_{h}(\cdot)(\cdot)$ and make it clear when properties of the specific problem (82) are used. With the Lagrangian $\mathcal{L}_{h}(u,z)=J(u)-A_{h}(u)(z)$ we formulate the error identity following the usual pathway (see for instance the proof of Theorem 1)
$J(u)-J(u_{h})=\frac{1}{2}\big{(}\mathcal{L}_{h}^{\prime}(u,z)(u-u_{h},z-z_{h})+\mathcal{L}_{h}^{\prime}(u_{h},z_{h})(u-u_{h},z-z_{h})\big{)}-A_{h}(u_{h})(z_{h})+{\cal R}^{(3)}.$ (86)
The very first term however will not vanish due to the non-conformity as $u_{h}$ and $z_{h}$ are not necessarily zero on the complete boundary $\partial\Omega$. As $L^{\prime}_{h}(u,z)(u,z)=0$ what remains is
$\mathcal{L}_{h}^{\prime}(u,z)(u-u_{h},z-z_{h})=-\mathcal{L}_{h}^{\prime}(u,z)(u_{h},z_{h}).$ (87)
For now, we leave this term and discuss it a little further down using problem (82) as an example. The second term or (86) is also not the standard weighted residual, as the adjoint variational form (85) is not the linearized adjoint of (83). Instead, we write
$\mathcal{L}^{\prime}_{h}(u_{h},z_{h})(u-u_{h},z-z_{h})$ $=-A_{h}(u_{h},z-z_{h})+J^{\prime}(u_{h})(u-u_{h})-A^{*}_{h}(u_{h})(u-u_{h},z_{h})$
$\quad+[A_{h}^{*}-A_{h}^{\prime}](u_{h})(u-u_{h},z_{h}),$ (88)
where $[A_{h}^{*}-A_{h}^{\prime}](u)(v,z):=A_{h}^{*}(u)(v,z)-A_{h}^{\prime}(u)(v,z)$ is the adjoint consistency error.
If we localize by introducing interpolations we must again take into account the non-conformity and finally reach the identity
$J(u)-J(u_{h})=\frac{1}{2}\rho_{h}(u_{h})(z-I_{h}z)+\frac{1}{2}\rho_{h}^{*}(u_{h},z_{h})(u-I_{h}u)+\rho_{h}(u_{h})(z_{h})+{\cal R}^{(3)}\\ -\frac{1}{2}\mathcal{L}^{\prime}_{h}(u,z)(u_{h},z_{h})+\frac{1}{2}[A_{h}^{*}-A_{h}^{\prime}](u_{h})(u-u_{h},z_{h})+\frac{1}{2}\mathcal{L}_{h}^{\prime}(u_{h},z_{h})(I_{h}u-u_{h},I_{h}z-z_{h})$ (89)
Here, $\rho_{h}(\cdot)(\cdot)$ is the usual residual and $\rho_{h}^{*}(u,z)(v)=J^{\prime}(u)(v)-A_{h}^{*}(u)(v,z)$. The first lines contains the usual terms: weighted primal and dual residuals, the iteration error and, finally, the third-order remainder. The second line consists of three parts: a term measuring the non-conformity of Nitsche’s method, a separate adjoint non-conformity term and, finally, another iteration error. In comparison to $\rho_{h}(u_{h})(z_{h})$, the term
$\mathcal{L}^{\prime}_{h}(u_{h},z_{h})(I_{h}u-u_{h},I_{h}z-z_{h})=J^{\prime}(u_{h})(I_{h}u-u_{h},z_{h})-A_{h}^{\prime}(u_{h})(I_{h}u-u_{h},z_{h})-A_{h}(u_{h})(I_{h}z-z_{h})$ (90)
is negligible as it is of higher order. It is an iteration error, however, it is additionally weighted with the errors $I_{h}u-u_{h}$ and $I_{h}z-z_{h}$.
Corollary 2 (Error estimation for Nitsche’s method applied to the model problem).
Let $u,z\in U:=H^{1}_{0}(\Omega)$ be solutions to (82) and (84), $u_{h},z_{h}\in U_{h}\subset H^{1}(\Omega)$ be approximations given by Nitsche’s method (83) and (85). Then, it holds
$J(u)-J(u_{h})\approx\frac{1}{2}\rho_{h}(u_{h})(z-I_{h}z)+\frac{1}{2}\rho_{h}^{*}(u_{h},z_{h})(u-I_{h}u)+\rho_{h}(u_{h})(z_{h})+\langle\partial_{n}z_{h},u_{h}\rangle_{\partial\Omega}.$ (91)
Proof.
We omit the third order remainder and discuss the three terms in the second line of (89). For the first non-conformity term we get with integration by parts and using $u=z=0$ on $\partial\Omega$ that
$\mathcal{L}^{\prime}_{h}(u,z)(u_{h},z_{h})$ $=J^{\prime}(u)(u_{h})-(\nabla z,\nabla u_{h})-(2uz,u_{h})+\langle u_{h},z\rangle_{\partial\Omega}-\frac{\gamma}{h}\langle u_{h},z\rangle_{\partial\Omega}$ (92)
$\quad+(f,z_{h})-(\nabla u,\nabla z_{h})-(u^{2},z_{h})+\langle\partial_{n}u,z_{h}\rangle_{\partial\Omega}-\frac{\gamma}{h}\langle u,z_{h}\rangle_{\partial\Omega}$
$=J^{\prime}(u)(u_{h})-(-\Delta z+2uz,u_{h})-\langle\partial_{n}z,u_{h}\rangle_{\partial\Omega}$
$\quad+(f,z_{h})-(-\Delta u+u^{2},z_{h})-\langle\partial_{n}u,z_{h}\rangle_{\partial\Omega}+\langle\partial_{n}u,z_{h}\rangle_{\partial\Omega}=-\langle\partial_{n}z,u_{h}\rangle_{\partial\Omega}.$
As $z\in H^{1}_{0}(\Omega)$ is not available, we introduce ±$\pm z_{h}$ and neglect the remainder term $\langle\partial_{n}(z-z_{h}),u_{h}\rangle_{\partial\Omega}$ as it carries an additional order of convergence over the remaining term $\langle\partial_{n}z_{h},u_{h}\rangle_{\partial\Omega}$.
For the second last term of (89) it holds using $u=z=0$ on $\partial\Omega$ that
$[A_{h}^{*}-A_{h}^{\prime}](u_{h})(u-u_{h},z_{h})$ $=-\langle\partial_{n}z_{h},u-u_{h}\rangle_{\partial\Omega}+\langle\partial_{n}(u-u_{h}),z_{h}\rangle_{\partial\Omega}$
$=\langle\partial_{n}z_{h},u_{h}\rangle_{\partial\Omega}+\langle\partial_{n}(u-u_{h}),z_{h}\rangle_{\partial\Omega}.$ (93)
We will neglect the second term, as it is qualitatively the same as (92) but carries the additional error $\partial_{n}(u-u_{h})$. Combining the above gives the postulated error approximation (91) ∎
The last term in (91) can be localized and attributed to the error indicators, e.g. by using the partition of unity method from Section 3.10. An alternative interpretation of this term however is to measure the violation of the boundary condition $u_{h}=0$ and it can be taken in its entirety to control the penalization parameter $\gamma>0$.